OpenStudy (anonymous):

Hi I got some questions wrong on an assignment i was wondering why I got them wrong and what was the correct answer may you help me?

OpenStudy (anonymous):

A video game club charges a fixed annual membership fee of $19 and $5 per video game rented. Let f(n) represent the total annual cost of renting n video games. Which of the following functions best represents the relationship between f(n) and n if the membership was decreased by $10 the next year? f(n) = 19n + 5 f(n) = 5n + 9 f(n) = 4n + 29 f(n) = 5n + 29 Points earned on this question: 0

OpenStudy (anonymous):

Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m) Points earned on this question: 0

OpenStudy (anonymous):

for the 1st one i got d and the second i got b i have others but i should just start w. those :)

OpenStudy (anonymous):

@hartnn @satellite73 @sammixboo

OpenStudy (ribhu):

the 1st seem to be b as the membership is decreased by $10 @hello_hey

hartnn (hartnn):

*decreased by 10* 19 -10 = ...

OpenStudy (anonymous):

9

hartnn (hartnn):

yep, so 5n +9

OpenStudy (anonymous):

may you explain everything to me like just pretend i was asking the question how would you explain it to me ?

OpenStudy (ribhu):

second one is y = f(m)

hartnn (hartnn):

sure, $5 per video game, so how many dollars for 'n' videogames?

OpenStudy (anonymous):

actually i get it thanks :)